Find the hypotenuse of each isosceles right triangle when the legs are of the given measure.
Given = 8
hint: answer _√2
(find the blank part of the answer)

Answer by richwmiller(17219)

2*8^2=c^2
2*64=c^2
128=c^2
c=8 sqrt(2)